# The modelling of the action potentials in myelinated nerve fibres

### v2

@@ Abstract
The classical Hodgkin-Huxley equations describe the propagation of an axon potential (AP) in unmyelinated axons. In many cases the axons have a myelin sheath and the experimental studies have then revealed significant changes in the velocity of APs. In this paper, a theoretical model is proposed describing the AP propagation in myelinated axons. As far as the velocity of an AP is affected, the basis of the model is taken after Lieberstein, who included the possible effect of inductance that might influence velocity, into the governing equation. The proposed model includes the structural properties of the myelin sheath: the -ratio (the ratio of the length of the myelin sheath and the Ranvier node) and g-ratio (the ratio of the inner-to-outer diameter of a myelinated axon) through parameter . It is demonstrated that the Lieberstein model can describe all the essential effects characteristic to the formation and propagation of an AP in an unmyelinated axon. Then a phenomenological model (a wave-type equation) for a myelinated axon is described including the influence of the structural properties of the myelin sheath and the radius of an axon. The numerical simulation using the physical variables demonstrates the changes in the velocity of an AP as well as the changes in its profile. These results match well the known effects from experimental studies.
@@
action potential, nerve fibre, myelinated axon, velocity, mathematical modelling1 Introduction
The celebrated Hodgkin-Huxley (HH) equations can describe the action potential (AP) in an unmyelinated axon taking into account sodium and potassium ion currents [1]. The strength of the HH equations is in the detailed description of the physics of ion currents but the cable equation describing the propagation of an AP is simplified by neglecting the inductance. There are several ways to improve this classical model towards a better description of the structural properties of axons and taking into account the accompanying effects. Many studies are devoted to the description of mechanical and/or thermal effects accompanying the propagation of an AP [2, 3, 4, 5] resulting in the formation of an ensemble of waves. These theoretical models are based on experimental results [6, 7, 8, 9, 10, 11]. Another important avenue of studies is related to modelling the behaviour of an AP in myelinated axons. The starting point for the modelling of an unmyelinated is a cylindrical tube embedded into the extracellular fluid. The barrier between extra- and intracellular fluids is a lipid bilayer with proteins. Such a biomembrane is sometimes modelled like a simple lipid bi-layer only [2]. In the case of the myelinated axon, this barrier has a myelin sheath which consists of multiple layers of a glial membrane composed of lipids and proteins and serves as an insulator [12, 13] or in some scenarios even as signal modulating element [14]. It means that under the myelin sheath the ion currents through the basic biomembrane (barrier) proposed by Hodgkin and Huxley [1], do not exist. However, the myelin sheath is interrupted by Ranvier nodes where the usual HH equations work. It is proposed that under the myelin sheath, the passive cable equation (the diffusion-type equation) describes the process and in Ranvier nodes, the usual HH equations work [15, 16]. However, it is demonstrated experimentally that neurons which have myelin sheathing often show higher AP velocities compared to similar unmyelinated neurons. This effect was already reported by Lillie in 1924 [17] and is nowadays referred to as the Lillie transition [18] or saltatory conduction (see [19]).
In this paper, a phenomenological model is proposed for describing the propagation of an AP in a myelinated axon. The most important question is how to model the possible changes in the propagation velocity of an AP due to the myelin sheath. The process is nonlinear and we follow a recommendation of Whitham [20]: “… one should not always turn too quickly to a search for the .” In the context of our equation, it means that we should keep the neglected inductance as small as it is (which we will see in Section 2.3 is about 40 [mH cm]). This idea is supported by Wang et al [21] who have argued that inductance is “a missing piece of neuroscience”. They state that the main source of the inductance is the myelin sheath. In general terms, however, inductance means that “there is a kind of biological structure that can store energy in a non-electrical form" [21]. If we keep the inductance in the cable equation then under the myelin sheath the propagation is described not by a diffusion (see [16]) equation but by a hyperbolic equation. This hypothesis may better explain the changes in the velocity of an AP. Given reports [22] noting the dependence of velocity on the diameter of a fibre, this relationship should be taken into account.
@@
Section 2 is devoted to the analysis of the Lieberstein [23] model for the unmyelinated axon. The results confirm the earlier analysis by Kaplan and Trujillo [28] but more details are given. As novel results, the profiles of ion currents and changes of the phenomenological variables , , and [1] during the propagation of the AP are also presented. It is shown that the head-on-collision of two APs leads to annihilation. The influence of the refractory period for the formation of consecutive APs is also analysed as well as the dependence of the velocity on the diameter of the axon. These well-known effects demonstrate the validity of the Lieberstein equation and encourage us to use this equation for the modelling of the APs in myelinated axons. Section 3 describes experimental results on myelinated axons focused on velocity changes [29, 30, 31], This analysis permits us to propose in Section 4 a phenomenological model describing the propagation of an AP in a myelinated axon influenced strongly by the structure of a myelin sheath (-ratio). The structure of the governing equations follows the Lieberstein model (see Section 2) where the inductance leads to a wave-like behaviour but the final velocity of an AP depends on ion currents. The numerical simulations demonstrate clearly the changes in the velocity under the myelin sheath of the propagating AP. Based on these results, a short discussion of forms of energy in an AP is presented in Section 4.2. Finally, in Section 5 the discussion is presented with conclusions. As far as the numerical simulations are carried on with physical variables (i.e., in physiologically viable range), the results could be better checked in experimental studies.
@@ 2 Lieberstein model for AP propagation on unmyelinated axon
Experimental studies of processes in the myelinated axons have demonstrated that the signal propagates faster than in unmyelinated axons [32, 33, 34]. Note that Cohen et al. [35] argue about elusiveness of the role of submyelin conduction. Although there are several explanations of this phenomenon [36, 26, 15] there is a need for constructing a mathematical model which allows to calculate of an AP profile in a myelinated axon with the velocities depending on structural properties of the myelin sheath. Consequently, in deriving a mathematical model, all possible phenomena that could influence the propagation velocity should be taken into account, even if some of these could be considered much smaller than others at first glance (see, for example, Whitham [20]).
@@ 2.1 Basic model
The cable equation used by Hodgkin and Huxley [1] for deriving the equations for the electrical signal propagation is parabolic, i.e., inductance is neglected. Indeed, its influence is small but following Whitham [20] we use here the full model of Lieberstein [23] derived directly from Maxwell equations [25]. This model is hyperbolic but the final velocity of the signal is influenced by ion currents like in the HH equations [1].
Here we go briefly over the key points from [23] demonstrate the importance of a threshold and annihilation and present the profiles of ion currents and phenomenological variables , , (not presented in [23]). We also add some comments to make it easier to follow later in the paper what is changed and why is changed when including the effects of myelination in the next Section.
@@
It is noted that the membrane current density (which is what is used in the HH equations [1]) is
@@
The structure and the physical interpretation of the governing equations (1) and (2) is easier to see if writing it as a single second-order PDE (see eq. (3) in [23]):
where on the left-hand side (LHS) there is a classical wave equation type operator while on the right-hand side (RHS) there are dissipation-type operators (first order partial time derivatives) which either remove energy from the system or add it to the system depending on the sign of the operator. In the context of signal propagation in nerve fibres, the ionic currents across the membrane change the membrane potential in a given location while some of that energy is lost during the diffusion-type propagation along the axon to the axon resistance. At the same time, the LHS operator ensures that part of the signal is propagating like a classical wave.
Returning to the system as two coupled eqs. (1) and (2) the membrane current is taken (this is what is different from regular telegraph equation) as it was proposed in the HH equations [1] as
where is the total membrane current density (positive direction is into the axon), is the membrane potential relative to the resting potential (depolarisation is negative), is the membrane capacity per unit area while and are currents which are defined through “equilibrium potentials" for Na and K ions while is “leakage current" representing all other ionic currents (Hodgkin and Huxley point mostly at chlorine Cl, however, Ca is also potentially relevant which some authors claim to be about as important as Na and K ions) and is chosen so that at resting potential leakage current across the membrane would be zero.
@@
The dimensions of coefficients in eqs. (8) and (9) are: , , , are dimensionless, ( denotes SI-unit: [Siemens]) while the membrane potential and the line axon current . As usual, the space dimension is , time dimension is , while in addition dimension is and dimension is .
The difference in keeping the inductance contrary to the celebrated classical HH equations [1] is the following. One possible way of making sense of the differences is by looking at the equations. In the case of the system of PDEs (eqs. (1) and (2) in [23]), it can be seen that there is a term in eq. (2). This can be interpreted as the rate of change of current along an axon. In the HH equations, the axon current is “hidden” in term . Another possibility is to look at the Lieberstein model in the form of one PDE (eqs. (3) and (6) in [23]). It can be seen in eq. (3) that two additional terms appear in the Lieberstein model compared to the classical HH equations – and . The second partial derivative is what makes the model hyperbolic and as pointed out by Kaplan and Trujillo [28], it influences the maximum velocity of the AP. However, the term is an even more notable difference as the time derivative of the ion current is not present in the classical HH equations. Observing eq. 6 in [28] it can be seen that inductance affects the ion currents and also time derivatives of gating variables arise. One should note that Kaplan and Trujillo [28] have calculated axon inductance from the movement of ions as if AP velocity is at µ and Cole and Baker have also investigated question of inductance in this context [37, 38]. It can be noted that similar models involving inductivity are in use in cardiophysiology, see, for example, [39].
@@ 2.2 Moving frame of reference and inductivity
Lieberstein [23] assumes that accounting for the inductivity (unlike in the HH equations where it was neglected) he can go into a moving frame of reference in a standard way introducing However, we should note that it could be, actually , where is velocity, as taking only a single direction means discarding the wave propagating in the opposite direction. It should be noted that using a moving frame of reference allows one to also derive an evolution equation (a single wave) for the nerve pulse [40]. Lieberstein [23] defined the velocity of the AP in an unmyelinated axon through axon radius , inductance and capacity as
@@
More recently, Fukasawa and Takizawa [41] have also investigated the question of electrical signal propagation velocity in an axon if inductivity is taken into account including the question of how to estimate the inductance parameter value.
In the present paper, we have opted to avoid going into a moving frame of reference preferring to solve the model as a coupled pair of PDEs and to keep the influence of for the sake of completeness, even if its influence on the behaviour of the solutions is arguably small.
2.3 Solutions of Lieberstein model
A numerical example (see Fig. 1) using pseudospectral method (PSM) (see Appendix A in [2]) and periodic boundary conditions demonstrating the evolution of solutions for model equations (8) and (9) is now presented. Briefly, the main point of the pseudospectral method is that the discrete Fourier transform (DFT) based (PSM) (see also [42]) can be used to represent variable in the Fourier space as

### v3

@@ Abstract
The initial version of the planned paper has gone through a major revision in 2025. First, the paper ended up growing a bit too long, and as a result of that, we decided to split it into two parts. The first part focuses on the model for the unmyelinated case and its behaviour, and the second part focuses on including the influence of myelination into the model. Second, when the initial version of the manuscript was going through the review process, it became evident that the way the content was presented was somewhat confusing for readers with a background in the experimental side of research into nerve processes. As a result, we went through a major revision, redoing all the numerical simulations with parameters that are closer to what Hodgkin and Huxley used in their classical paper from 1952, where the Hodgkin-Huxley model was initially introduced. The second major change was to change the logic how the specific inductance value is chosen for the numerical example - in the previous version it was chosen by aiming for a specific propagation velocity when the axon radius was chosen as 1 micrometre, in the updated version the value is chosen to get the AP propagation velocity which was experimentally observed in the HH 1952 paper at the same parameters as were used in that paper.
Part 1 - On hyperbolicity for nerve pulse propagation in axons.
The classical Hodgkin-Huxley (HH) model describes the propagation of an axon potential (AP) in unmyelinated axons. The hypothesis is that the AP propagation in axons depends not only on the HH ion mechanism but also on capacitance and inductance.
In this paper, we revisit a model proposed by Lieberstein for describing propagating AP in unmyelinated axon, including the possible effect of inductance that might influence velocity, into the governing equation. In many cases the axons have a myelin sheath and the experimental studies have then revealed significant changes in the velocity of APs. Next, the goal is to modify Lieberstein model further to include the influence of myelination on the AP dynamical behaviour. However, before we can do that we have to check that the solutions of the governing equations fulfil all the essential requirements for describing the nerve signalling in a physiologically plausible way.
The numerical simulation using the physical variables demonstrates the changes in the velocity of an AP as well as the changes in its profile. These results match well the known effects from experimental studies.
Part 2 - The modelling of the action potentials in myelinated nerve fibres.
The classical Hodgkin-Huxley model describes the propagation of an axon potential (AP) in unmyelinated axons. In many cases the axons have a myelin sheath and the experimental studies have then revealed significant changes in the velocity of APs. In this paper, a theoretical model is proposed describing the AP propagation in myelinated axons. As far as the velocity of an AP is affected, the basis of the model is taken after Lieberstein, who included the possible effect of inductance that might influence velocity, into the governing equation. The proposed model includes the structural properties of the myelin sheath: the -ratio (the ratio of the length of the myelin sheath and the Ranvier node) and g-ratio (the ratio of the inner-to-outer diameter of a myelinated axon) through parameter . The Lieberstein model can describe all the essential effects characteristic to the formation and propagation of an AP in an unmyelinated axon. Then a phenomenological model (a wave-type equation) for a myelinated axon is described including the influence of the structural properties of the myelin sheath and the radius of an axon. The numerical simulation using the physical variables demonstrates the changes in the velocity of an AP. These results match well the known effects from experimental studies.
@@
[…]
Part 1 – On hyperbolicity for nerve pulse propagation in axons
1 Introduction
The celebrated Hodgkin-Huxley (HH) model describes the dynamics of an action potential (AP) in unmyelinated axons [23]. This model explicitly describes the role of ion currents in forming an asymmetric AP in an unmyelinated nerve fibre. In deriving the governing equations for a squid giant axon, Hodgkin and Huxley neglected the role of inductance and their model is based on the cable equation which is a diffusion-type equation [45]. Their model (in the form of an ordinary differential equation (ODE)) was useful for calculating an AP at a certain space point in time. In order to get a propagating wave, they made a few additional assumptions to get back to governing equations that are in the form of partial differential equations (PDE). It should be noted that partly the reasoning behind preferring the ODE form of HH model was the performance of computational resources back in 1950’ies. Clearly, the computational resources have increased by many orders of magnitude over the past 75 years. Alternatively, based on part of the work done by Hodgkin and Huxley [23], Lieberstein [31] proposed a model that allows a propagating AP signal by opting to keep the inductivity in a way that does not require these additional assumptions.
1.1 Lieberstein hypothesis
Lieberstein [31] hypothesis – the AP propagation in axons depends not only on the HH ion mechanism but also on capacitance and inductance.
1.2 The question of inductance in the context of nerve fibres
The discussion of the role of inductance started even before the derivation of the HH model. Cole has analysed the influence of inductance in the process of forming signals in axons [5, 6] and concluded that it may be worth including inductance in the model. Lieberstein has modified the HH model by returning to basics and kept inductance in the model [31] and added the mechanism of ion currents proposed by Hodgkin and Huxley [23]. In this model, the governing equation is hyperbolic and describes a propagating wave. Having a hyperbolic governing equation (wave-equation type) is beneficial as it ensures that the process remains causal (i.e., velocity is finite) as opposed to parabolic (heat-equation type) where the signal can have infinite propagation velocity (i.e., breaking causality). Recently, Wang et al [57] have stressed the importance of inductance in processes of myelinated axons. In this context, the Lieberstein model may be more informative describing the process in more detail. The original governing equations of Lieberstein [31] were solved by the finite difference method and the results obtained by numerical simulation demonstrated some wave profiles in time. Below we shall explain the derivation of the Lieberstein governing equations and demonstrate that the model can describe all the essential effects characteristic to the formation and propagation of an AP in unmyelinated axons. In addition, the profiles of ion currents and phenomenological variables proposed by Hodgkin and Huxley [23] are analysed based on the solution of the Lieberstein equation. In this way, the results of this study complete the analysis of the Lieberstein model preparing the ground for the modelling of the AP in the myelinated axons.
1.3 Structure of the paper
@@
We start with a brief overview of the structure of nerve cells in Section 2 and a brief overview of classical models in such a context in Section 3. Section 4 is devoted to the analysis of the Lieberstein [31] model for the unmyelinated axon including the brief overview of assumptions made when deriving the noted model. Section 5 contains the profiles of an AP found by the numerical simulation using the physical units [23, 31] demonstrating the accuracy of the model. The obtained results confirm the earlier analysis by Kaplan and Trujillo [29] but more details are given. The profiles of ion currents and changes of the phenomenological variables , , and [23] during the propagation of the AP are also presented. It is shown that the head-on-collision of two APs leads to annihilation. The influence of the refractory period for the formation of consecutive APs is also analysed as well as the dependence of the velocity on the diameter of the axon. These well-known effects demonstrate the validity of the Lieberstein model and encourage us to use this equation for the modelling of the APs. Finally, in Section 6 the discussion is presented with conclusions. As far as the numerical simulations are carried out with physical variables (i.e., in physiologically viable range), the results could be better checked in experimental studies.
2 Brief description of the physical structure of an axon
The main structural element in neural networks is the axon along which an electrical signal (the action potential (AP)) propagates from the cell body to the nerve terminal. An axon can be modelled as a tube in a certain environment [25]. Inside the tube is the axoplasmic fluid (shortly axoplasm), and the wall (in terms of continuum mechanics) of the tube has a bi-layered lipid structure called biomembrane. The biomembrane is composed of two layers of amphiphilic phospholipids and also includes proteins which are responsible for forming the ion channels. Such a simple biomembrane is called unmyelinated but in many cases, the biomembrane has an additional myelin sheath composed of multiple layers of a glial membrane. In this case, the axon is called myelinated. The sheath is interrupted by nodes of Ranvier where ion channels are active like in the case of unmyelinated axon. Under the myelin sheath, typically there are no active ion channels. The idealised schemes for a nerve cell and unmyelinated axon are shown in Fig. 1.
3 Modelling
The strength of the HH model is in the detailed description of the mechanism of ion currents but the cable equation describing the propagation of an AP is simplified by neglecting the inductance. There are several ways to improve this classical model towards a better description of the structural properties of axons and taking into account the accompanying effects. Many studies are devoted to the description of mechanical and/or thermal effects accompanying the propagation of an AP [11, 9, 4, 28] resulting in the formation of an ensemble of waves. These theoretical models are based on experimental results [26, 51, 54, 59, 55, 52].
The starting point for the modelling of an unmyelinated axon is a cylindrical tube embedded into the extracellular fluid. The barrier between extra- and intracellular fluids is a lipid bilayer with proteins. Such a biomembrane is sometimes modelled like a simple lipid bi-layer only [11]. In the case of the myelinated axon, this barrier has a myelin sheath which consists of multiple layers of a glial membrane composed of lipids and proteins and serves as an insulator [7] or in some scenarios even as signal modulating element [56, 14]. It means that under the myelin sheath the ion currents through the basic biomembrane (barrier) proposed by Hodgkin and Huxley [23], do not exist. However, the myelin sheath is interrupted by Ranvier nodes where the usual HH model is adequate. It is proposed that under the myelin sheath, the passive cable equation (the diffusion-type equation) describes the process and in Ranvier nodes, the classical HH model can be used [15, 20]. However, it is demonstrated experimentally that neurons which have myelin sheathing often show higher AP velocities compared to similar unmyelinated neurons. This effect was already reported by Lillie in 1924 [32] and is nowadays referred to as the Lillie transition [60] or saltatory conduction (see [25]). Before including the effect of myelination on the AP propagation we have to map the behaviour of the model in the simpler unmyelinated case.
In this paper, a model based on Lieberstein [31] is studied which describes the propagation of an AP in an axon. This is the first step, we map the behaviour of the solutions in the unmyelinated case before we move to modify the model for the myelinated case in the future. The model is nonlinear and we follow a recommendation of Whitham [58]: “… one should not always turn too quickly to a search for the ” In the context of this model, it means that we should keep the neglected inductance as small as it is. This idea is supported by Wang et al [57] who have argued that inductance is “a missing piece of neuroscience”. In general terms, however, inductance means that “there is a kind of biological structure that can store energy in a non-electrical form” [57]. If we keep the inductance in the cable equation then the propagation is described not by a diffusion (see [20]) equation but by a hyperbolic equation. This hypothesis may better explain the changes in the velocity of an AP. Noting the dependence of velocity on the diameter of a fibre [42] in the earlier studies and this relationship should also be taken into account.
In what follows, the model of Lieberstein [31] is a basic one to describe the AP propagation in axons. This model includes the HH ion currents but in addition, includes the inductance as it follows from the Maxwell equations [34]. However, before moving on to the cases with higher complexity (i.e., including effects of myelination in the future) we should make sure that the basic model captures all the essential effects of the AP propagation.
@@ 4 Lieberstein model for AP propagation on unmyelinated axon
The goal is to model propagation of AP along axon which is a component of the nerve cell located between the cell body and synapse (where the signal is transmitted to the next nerve cell, see also Fig. 1). Lieberstein model can be considered as a modification of the HH model as underlying major assumptions during the derivation are the same, like, for example, description of the dynamics of key protein channels (ion channels). Here we shall demonstrate that the Lieberstein model is able to describe all the important phenomena in unmyelinated axons and will be a great canditate for further modifications for the purpose of modelling AP propagation in myelinated axons.
@@ 4.1 Basic model
The cable equation [15] used by Hodgkin and Huxley [23] for deriving the equations for the electrical signal propagation is parabolic, i.e., inductance is neglected. Indeed, its influence is small but following Whitham [58] we use here the full model of Lieberstein [31] derived directly from Maxwell equations [34]. This model is hyperbolic but the final velocity of the signal is influenced by ion currents like in the HH equations [23] as noted before.
Here we go briefly over the key points from [31] demonstrating the importance of a threshold and annihilation and presenting the profiles of ion currents and phenomenological variables , , (not presented in [31]). We add some comments related to the modifications we plan to do in the future when including the effect of myelination on the AP propagation along the axon.
@@
It is noted that the membrane current density (which is what is used in the HH model [23]) is
@@
It is easier to understand the structure and the physical interpretation of governing equations (1) and (2) by transferring the system of equations to a single second-order PDE (see eq. (3) in [31]):
where on the left-hand side (LHS) is a classical wave equation type operator while on the right-hand side (RHS) are dissipation-type operators (first order partial time derivatives). These operators either remove energy from the system or add it to the system depending on the sign of the operator. In the context of signal propagation in nerve fibres, the ionic currents across the membrane change the membrane potential in a given location while some of that energy is lost during the diffusion-type propagation along the axon to the axon resistance. At the same time, the LHS operator ensures that part of the signal is propagating like a classical wave.
Returning to the system as two coupled eqs. (1) and (2), the membrane current is taken (this is what is different from regular telegraph equation) as it was proposed in the HH model [23] as
where is the total membrane current density (positive direction is into the axon), is the membrane potential relative to the resting potential (depolarisation is negative), is the membrane capacity per unit area while and are currents which are defined through “equilibrium potentials” for Na and K ions while is “leakage current” representing all other ionic currents (Hodgkin and Huxley refer to chlorine Cl, however, Ca is also relevant for some excitable cells) and is chosen so that at resting potential leakage current across the membrane would be zero.
@@
Equations (8) and (9) can be solved numerically with the pseudospectral method. It should be noted that Lieberstein [31] remarks that as is small then , and the influence of is typically assumed to be negligible. We take in the following numerical example. We performed also a few numerical simulations with (just taking it one order of magnitude smaller than to check if it does anything significant if taken large enough) but observed its influence to be small on the behaviour of the numerical solutions.
The difference in keeping the inductance contrary to the celebrated classical HH model [23] is the following. One possible way of making sense of the differences is by looking at the equations. In the case of the system of PDEs (eqs. (1) and (2) in [31]), it can be seen that there is a term in eq. (2). This can be interpreted as the rate of change of current along an axon. In the HH equations, the axon current is “hidden” in term . Another possibility is to look at the Lieberstein model in the form of one PDE (eqs. (3) and (6) in [31]). It can be seen in eq. (3) that two additional terms appear in the Lieberstein model compared to the classical HH equations – and . The second partial derivative is what makes the model hyperbolic and as pointed out by Kaplan and Trujillo [29], it influences the maximum velocity of the AP. While the term is another notable difference as the time derivative of the ion current is not present in the classical HH model. Observing eq. (6) in [29] it can be seen that inductance affects the ion currents and also time derivatives of gating variables arise. One should note that Kaplan and Trujillo [29] have calculated axon inductance from the movement of ions as if AP velocity is at µ and Cole and Baker have also investigated question of inductance in this context [5, 6]. It can be noted that similar models involving inductivity are in use in cardiophysiology, see, for example, [41].
@@ 4.2 Moving frame of reference and inductivity
Lieberstein [31] assumes that accounting for the inductivity (unlike in the HH equations where it was neglected), he can go into a moving frame of reference in a standard way introducing However, we should note that it could be, actually , where is velocity, as taking only a single direction means discarding the wave propagating in the opposite direction. We remark that using a moving frame of reference allows one to also derive an evolution equation (a single wave) for the nerve pulse [10]. Lieberstein [31] defined the velocity of the AP in an unmyelinated axon through axon radius , inductance and capacity as
@@
More recently, Fukasawa and Takizawa [19] have also investigated the problem of electrical signal propagation velocity in an axon if inductivity is taken into account including the question of how to estimate the inductance parameter value.
In the present paper, we have opted to avoid going into a moving frame of reference preferring to solve the model as a coupled pair of PDEs and to keep the influence of for the sake of completeness, even if its influence on the behaviour of the solutions is arguably small. We comment that first, for our chosen numerical solving method it is more convenient to solve a coupled pair of PDEs instead of a single higher order PDE with mixed partial derivatives and second, the model will be easier to modify in the future to include effects of myelination in the paired PDE form where the action potential across the membrane and ionic currents along the axon are separately represented.
5 Solutions of Lieberstein model
The system is solved using the pseudospectral method (PSM) (see Appendix A in [11] for details) and periodic boundary conditions demonstrating the evolution of solutions for model equations (8) and (9) (see Fig. 2). For initial condition we generate a narrow localised pulse for AP in the middle of the 1D space domain at time as where where is amplitude of the pulse (-15 [mV] unless stated otherwise), is the width parameter of the pulse, is number of 2 sections in space. For axial current we take zero initial value and values for are given in Table 1. Briefly, the main point of the pseudospectral method is that the discrete Fourier transform (DFT) based (PSM) (see also [17]) can be used to represent variable in the Fourier space as